# Multimodal Adversarial Defense for Vision-Language Models by Leveraging One-To-Many Relationships

### v2

@@ 4.3 Leveraging one-to-many relationships for robustness generalization
While MAT addresses the multimodality of VL tasks, we hypothesize that robustness can be further improved by addressing also the one-to-many nature of image-text data. Thus, we propose Multimodal Augmented Adversarial Training (MAT+), which incorporates one-to-many augmentations to better approximate the true multimodal distribution of images and texts. Notably, this data-driven strategy for enhancing robustness is applicable to any VL model trained on image-text pairs.
@@ 4.3.1 Approximating the multimodal distribution by modeling ambiguity via augmentations
The true distribution of the data $\mathcal{D}^{\star}$ is inaccessible, but the dataset $\mathcal{\hat{D}}$ serves as an approximation, and its quality affects robustness significantly. In the field of image classification, Gowal et al. (2021) proved that adding image augmentations to the training data $\mathcal{\hat{D}}_{\text{tr}}$ produces a $\mathcal{\hat{D}}$ closer to $\mathcal{D}^{\star}$. This improves AT under the assumption of a deterministic closed-set image-to-label mapping. However, their theory does not consider the ambiguity of VL data, where a single image can have multiple valid descriptions and vice versa (Fig. 1). Thus, for VL data, multimodal AT with (1:1) image-text pairs only provides a limited $\mathcal{\hat{D}}$. To model this ambiguity, we generate diverse one-to-many (1:N) and many-to-one (N:1) image-text pairs through augmentations, improving the approximation of Eq. 5.
Modeling the ambiguity of text descriptions. Let us suppose a $\mathcal{\hat{D}}_{\text{tr}}$ consisting of deterministic (1:1) image-text pairs $(I,T)$. Relying solely on these pairs produces a weak $\mathcal{\hat{D}}$ because the original text description $T$ in $\mathcal{\hat{D}}_{\text{tr}}$ typically provides a partial or subjective depiction of the image $I$ (Fig. 1). In other words, $\mathcal{\hat{D}}_{\text{tr}}$ is typically constructed under the premise that $(I,T)$ are not ideal, but an approximation of the true image-to-text model $g_{\text{T}}^{\star}:\mathbb{R}^{d_{\text{I}}}\rightarrow\mathbb{R}^{d_{\text{T}}}$ (i.e., a “perfect” human annotator capable of generating exhaustive detailed descriptions). However, striving for a true $g_{\text{T}}^{\star}$ is neither desirable nor practical; language compresses visual information by focusing on human interests, naturally involving ambiguity, and approximating $g_{\text{T}}^{\star}$ deterministically would require a description detailed down to the individual pixel, which is linguistically unrealistic.
@@ 4.3.2 Conditions for effective augmentations
Not all $\Phi$ and $\Psi$ produce a model that can generate valid image-text pairs ×$\mathcal{I}\times\mathcal{T}\subseteq\mathbb{R}^{d_{\text{I}}}\times\mathbb{R}^{d_{\text{T}}}$, that is, paired data that effectively refines $\mathcal{\hat{D}}$ towards $\mathcal{D}^{\star}$. We can develop the theory of Gowal et al. (2021) for image classification to address multimodal VL augmentations. Let $p$ be the probability measure corresponding to $\mathcal{D}^{\star}$, for which every valid pair $(I,T)\in\mathcal{I}\times\mathcal{T}$ has non-zero probability: $p(I,T)>0$. Similarly, let $\hat{p}$ be the probability measure corresponding to $\mathcal{\hat{D}}$. The sufficient condition for $\mathcal{\hat{D}}$ to be an effective approximation is as follows:
@@
(Accurate approximated distribution) The approximated data distribution $\mathcal{\hat{D}}$ and true data distribution $\mathcal{D}^{\star}$ must be equivalent: $\forall(I,T)\in\mathcal{I}\times\mathcal{T},p(I,T)=\hat{p}(I,T).$
@@
(Distribution shift) Even if augmented image-text pairs are aligned, excessive augmentations cause distribution shifts. This also applies to the unimodal case (Gowal et al., 2021).
To make sure our augmentations are effective, we introduce three measurable properties as follows:
  • Distribution gap: The Fréchet distance (Fréchet, 1936) between the distributions of the original and augmented samples, $d_{\text{F}}(\mathcal{\hat{D}}_{\text{tr}},\mathcal{\hat{D}}_{\text{aug}})$, where $\mathcal{\hat{D}}_{\text{aug}}=\{(I_{\text{aug}},T),(I,T_{\text{aug}})\}^{N}_{i=1}$. This metric, widely used for evaluating generative models (e.g., FID (Heusel et al., 2017)), quantifies the distribution gap.
Our results in Section 6 show that cross-modal augmentations (image $\leftrightarrow$ text) achieve better alignment than intra-modal augmentations (image $\to$ image, text $\to$ text), which cannot capture image-text interactions. Moreover, achieving both sufficient diversity and minimal distribution gap is harder to achieve in image augmentations due to the high dimensionality and degree-of-freedom of the image space.
@@ 5 Experimental settings
Our evaluation is focused mainly on improving robustness for image text retrieval (ITR), the most fundamental VL task. For this, we apply MAT to three representative VL models: CLIP, ALBEF, and BLIP. Finally, we also evaluate additional VL tasks, i.e., ALBEF on visual grounding (VG) and BLIP on image captioning (IC).
Method Img aug. Text. aug. Clean Co-Attack SGA
TR@1 IR@1 TR@1 IR@1 TR@1 TR@5 IR@1 IR@5
Finetune 92.1 77.2 11.0 6.7 0.6 2.0 0.6 2.6
@@
I2T(div-Caps) 84.7 $\uparrow$1.0 69.2 $\uparrow$1.7 60.3 $\uparrow$3.4 39.6 $\uparrow$2.6 40.3 $\uparrow$2.8 66.4 $\uparrow$1.3 27.8 $\uparrow$3.0 54.3 $\uparrow$4.5
I2T(Human) 85.7 $\uparrow$2.0 71.9 $\uparrow$4.4 63.3 $\uparrow$6.4 44.0 $\uparrow$6.9 45.6 $\uparrow$8.1 70.4 $\uparrow$5.3 32.2 $\uparrow$7.4 58.4 $\uparrow$8.6
Basic(RandAug) 84.1 $\uparrow$0.4 67.1 $\downarrow$0.4 57.8 $\uparrow$0.9 37.3 $\uparrow$0.2 35.6 $\downarrow$1.9 64.3 $\downarrow$0.8 24.4 $\downarrow$0.4 50.2 $\uparrow$0.4
T2I(SD) 83.3 $\downarrow$0.4 68.4 $\uparrow$0.9 57.2 $\uparrow$0.3 37.8 $\uparrow$0.7 37.9 $\uparrow$0.4 63.4 $\downarrow$1.7 25.2 $\uparrow$0.4 50.1 $\uparrow$0.3
T-I2I(SD) 83.8 $\uparrow$0.1 69.1 $\uparrow$1.6 59.7 $\uparrow$2.8 38.0 $\uparrow$1.0 39.3 $\uparrow$1.8 66.1 $\uparrow$1.0 25.8 $\uparrow$1.0 51.2 $\uparrow$1.5
Table 1: Comparison of CLIP trained on the Flickr30k dataset for image-text retrieval (ITR) under the no-attack scenario (Clean) and multimodal attacks (Co-Attack, SGA), reporting R@k for text retrieval (TR@k) and image retrieval (IR@k).
Method Img aug. Text. aug. Clean Co-Attack SGA
TR@1 IR@1 TR@1 IR@1 TR@1 TR@5 IR@1 IR@5
Finetune 66.6 50.1 2.9 1.8 0.1 0.5 0.1 0.6
@@
I2T(div-Caps) 56.7$\uparrow$0.9 39.9$\downarrow$0.8 30.3$\downarrow$1.1 19.4$\downarrow$0.2 18.9$\uparrow$1.2 40.2$\uparrow$1.1 12.5$\uparrow$0.2 30.2$\downarrow$0.6
I2T(Human) 58.9$\uparrow$3.1 43.1$\uparrow$2.4 33.6$\uparrow$2.2 21.9$\uparrow$2.3 21.3$\uparrow$3.6 43.2$\uparrow$4.1 14.5$\uparrow$2.2 34.6$\uparrow$3.8
Basic(RandAug) 55.9$\uparrow$0.1 40.7$\uparrow$0.0 30.7$\downarrow$0.7 19.6$\uparrow$0.1 18.3$\uparrow$0.6 40.1$\uparrow$1.0 12.6$\uparrow$0.3 31.3$\uparrow$0.6
T2I(SD) 54.2$\downarrow$1.6 38.3$\downarrow$2.4 28.5$\downarrow$2.9 18.1$\downarrow$1.5 17.2$\downarrow$0.5 37.5$\downarrow$1.6 11.7$\downarrow$0.6 29.9$\downarrow$0.9
T-I2I(SD) 57.1$\uparrow$1.3 41.7$\uparrow$1.0 31.3$\downarrow$0.0 20.2 $\uparrow$0.7 18.8 $\uparrow$1.1 39.7 $\uparrow$0.6 12.6 $\uparrow$0.2 31.0 $\uparrow$0.3
Table 2: Comparison of CLIP trained on the COCO dataset for image-text retrieval (ITR) under the no-attack scenario (Clean) and multimodal attacks (Co-Attack, SGA), reporting R@k for text retrieval (TR@k) and image retrieval (IR@k).
Datasets. We use the commonly used Flickr30k (Plummer et al., 2015) and the COCO (Chen et al., 2015) datasets for ITR. While these datasets contain five captions per image, our baseline training uses 1:1 image-text pairs (taking the first annotated caption), reflecting the typical setup when fine-tuning with real-world data, such as data collected from the internet, where 1:1 pairings are more common. For VG, we use RefCOCO+ (Kazemzadeh et al., 2014), and for IC, we use COCO. Please see Appendix A for details.
Evaluation metrics. We use the recall@k (R@k) for evaluating ITR, including both image-to-text and text-to-image retrieval. For VG we use the accuracy in localizing the regions corresponding to the text descriptions, while IC evaluates the quality of generated captions using diverse metrics. Details are in Appendix A.
Adversarial attacks. We evaluate our defense methods against multimodal adversarial attacks, Co-Attack (Zhang et al., 2022) and SGA (Lu et al., 2023). The perturbation constraints are $\epsilon_{I}=2/255$ ($L_{\infty}$-norm) for image and one token for text, following (Zhang et al., 2022; Lu et al., 2023). Additional results on unimodal attacks, including PGD and BERT-Attack, are in Appendix C.
Baseline methods. Since there do not exist defense methods tailored for multimodal attacks, we compare our method with the unimodal AT methods closest to our setting proposed for CLIP:
  • FARE (Schlarmann et al., 2024): An unsupervised adversarial fine-tuning scheme for CLIP, which focuses on obtaining a robust CLIP vision encoder. Since it is unsupervised, we can apply it to our ITR setting as-is.
  • TeCoA-ITR: Since the original TeCoA (Mao et al., 2022) is supervised, we extend it to be used in ITR. While TeCoA fine-tunes the vision encoder for image classification, TeCoA-ITR fine-tunes both vision and text encoders using Eq. 7 to generate adversarial images.
Training details. We fine-tune the pre-trained CLIP-ViT-B/16, ALBEF-14M, and BLIP w/ ViT-B models using MAT. For this, we generate adversarial images using 2-step-PGD (perturbation size of 2/255 in $l_{\infty}$-norm), and adversarial texts using BERT-attack (perturbation size of 1 token). Please see the detailed hyperparameters in Appendix A.4.
@@
[…]
@@
[…]
@@
    • Intra-modal: EDA (Wei & Zou, 2019) for basic word-level edits. LLM-based rewriting (Fan et al., 2024) is shown in Appendix C.
    • Cross-modal: Image-to-text (I2T) generation using diverse prompts with InternVL (Chen et al., 2024) (“I2T(div-Caps)”), and human-generated captions from the remaining 4 annotations of Flickr30k and COCO (“I2T(Human)”).
@@
    • Intra-modal: RandAug (Cubuk et al., 2020), randomly applying affine transformations and color distortions.
    • Cross-modal: Stable Diffusion (SD) for text-to-image (T2I) (Rombach et al., 2022) and text-guided image-to-image (T-I2I) (Meng et al., 2021).
@@ 6.1 Image-text retrieval (ITR)
Table 1 and 2 present the results for ITR using CLIP trained on Flickr30k and COCO, respectively. Table 3 presents the results for ALBEF trained on the same datasets, while Tab. 4 evaluates BLIP on COCO.
@@ 6.1.1 Effectiveness of Multimodal Adversarial Training
By introducing multimodal adversarial perturbations, MAT consistently achieves significantly greater robustness against multimodal attacks than the unimodal AT methods, FARE and TeCoA-ITR, which focus solely on image perturbations during AT. The improvements are substantial and consistent for CLIP on Flickr30k (Tab. 1) and COCO (Tab. 2), as well as ALBEF on both datasets (Tab. 3). These results highlight the necessity of defense strategies tailored for multimodal perturbations.

### v3

@@ 4.3 Leveraging one-to-many relationships for robustness generalization
While MAT addresses the multimodality of VL tasks, the robustness can be further improved by leveraging also the one-to-many nature of image-text data. Thus, we propose Multimodal Augmented Adversarial Training (MAT+), which incorporates one-to-many augmentations to better approximate the true multimodal distribution of images and texts. Notably, this data-driven strategy for enhancing robustness is applicable to any VL model trained on image-text pairs.
@@ 4.3.1 Approximating the multimodal distribution by modeling ambiguity via augmentations
The true distribution of the data $\mathcal{D}^{\star}$ is inaccessible, but the dataset $\mathcal{\hat{D}}$ serves as an approximation, and its quality affects robustness significantly. In the field of image classification, gowal2021improving proved that adding image augmentations to the training data $\mathcal{\hat{D}}_{\text{tr}}$ produces a $\mathcal{\hat{D}}$ closer to $\mathcal{D}^{\star}$. This improves robustness under the assumption of a deterministic closed-set image-to-label mapping. However, their theory does not consider the ambiguity of VL data, where a single image can have multiple valid descriptions and vice versa (Fig. 1). Thus, for VL data, multimodal AT with (1:1) image-text pairs only provides a limited $\mathcal{\hat{D}}$. To model this ambiguity, we generate diverse one-to-many (1:N) and many-to-one (N:1) image-text pairs through augmentations, improving the approximation of Eq. 6.
Modeling the ambiguity of text descriptions. Let $\mathcal{\hat{D}}_{\text{tr}}$ be a dataset of deterministic (1:1) image-text pairs $(I,T)$. Relying solely on these pairs produces a weak $\mathcal{\hat{D}}$ because the original text $T$ often provides only a partial or subjective depiction of the image $I$ (Fig. 1). In other words, $\mathcal{\hat{D}}_{\text{tr}}$ approximates the true image-to-text mapping $g_{\text{T}}^{\star}:\mathbb{R}^{d_{\text{I}}}\rightarrow\mathbb{R}^{d_{\text{T}}}$, a “perfect” human annotator capable of generating detailed descriptions. However, striving for a true $g_{\text{T}}^{\star}$ is neither desirable nor practical; language compresses visual information by focusing on human interests, naturally involving ambiguity. Achieving $g_{\text{T}}^{\star}$ would require pixel-level descriptions, which are linguistically unrealistic.
@@ 4.3.2 Conditions for effective augmentations
Not all $\Phi$ and $\Psi$ produce a model that can generate valid image-text pairs $\mathcal{I}\times\mathcal{T}\subseteq\mathbb{R}^{d_{\text{I}}}\times\mathbb{R}^{d_{\text{T}}}$, that is, paired data that effectively refine $\mathcal{\hat{D}}$ toward $\mathcal{D}^{\star}$. We can develop the theory of gowal2021improving for image classification to address multimodal VL augmentations. Let $p$ be the probability measure corresponding to $\mathcal{D}^{\star}$, for which every valid pair $(I,T)\in\mathcal{I}\times\mathcal{T}$ has non-zero probability: $p(I,T)>0$. Similarly, let $\hat{p}$ be the probability measure corresponding to $\mathcal{\hat{D}}$. The sufficient condition for $\mathcal{\hat{D}}$ to be an effective approximation is as follows:
@@
(Accurate approximation) The approximated data distribution $\mathcal{\hat{D}}$ and true data distribution $\mathcal{D}^{\star}$ must be equivalent: $\forall(I,T)\in\mathcal{I}\times\mathcal{T},p(I,T)=\hat{p}(I,T).$
@@
(Distribution shift) Even if augmented image-text pairs are aligned, excessive augmentations cause distribution shifts. This also applies to the unimodal case [gowal2021improving].
Our results in Sec. 6 show that MAT+ with well-chosen augmentations avoiding these three cases significantly enhances robustness.
@@ 5 Experimental settings
We evaluate the multimodal robustness of defense methods on VLMs (i.e., CLIP and ALBEF) in their original tasks they were evaluated in, image-text retrieval (ITR) and visual grounding (VG). We also include BLIP, a VLM capable of text generation for image captioning (IC).
Method Img aug. Text aug. Flickr30k COCO
Clean SGA Clean SGA
TR@1 IR@1 TR@1 IR@1 TR@1 IR@1 TR@1 IR@1
Finetune 92.1 77.2 0.6 0.6 66.6 50.1 0.1 0.1
FARE 75.9 61.0 27.1 21.0 45.2 32.3 9.1 6.9
TeCoA-ITR 83.1 68.2 27.5 17.6 58.0 41.6 9.6 6.2
(ours) MAT 84.6 67.7 36.4 24.9 55.8 40.7 17.7 12.3
@@
[…]
Table 1: Comparison of CLIP trained on the Flickr30k and COCO datasets for image-text retrieval (ITR) under the no-attack scenario (Clean) and multimodal attack (SGA), reporting R@k for text retrieval (TR@k) and image retrieval (IR@k).
@@
(ours) MAT+ Basic(EDA) 82.2 $\uparrow$0.2 67.9 $\uparrow$1.6 44.6 $\downarrow$2.5 31.2 $\downarrow$1.7 63.9 $\uparrow$0.0 46.8 $\uparrow$0.6 31.5 $\uparrow$0.3 20.9 $\downarrow$0.3
@@
Adversarial Training Config. Clean PGD (image) BERT-Attack (text) SGA (multimodal) Time cost
Order
Image Attack
(Obj., Optim.)
Text Attack
(Obj., Optim.)
Trained
params.
IR@1 TR@1 IR@1 TR@1 IR@1 TR@1 IR@1 TR@1 (sec/iter) (relative)
Finetune - - - All 92.1 77.2 11.9 10.1 75.4 53.1 0.6 0.6 1.13 $\times$ 0.13
FARE I (Uni, PGD-10) - Vision 75.9 61.0 69.7 55.1 53.2 40.2 27.1 21.0 7.81 $\times$ 0.89
TeCoA-ITR I (Cross, PGD-10) - All 83.1 68.2 77.7 61.9 64.7 42.7 27.5 17.6 10.29 $\times$ 1.17
MAT T → I (Cross, PGD-2) (Cross, BERT) All 83.7 67.5 77.4 61.4 72.2 51.1 37.5 24.8 8.79 -
Objective ablation
(1-1) T → I (Uni, PGD-2) (Cross, BERT) All 90.6 76.5 70.7 57.0 80.4 59.4 16.2 11.6 8.45 $\times$ 0.96
(1-2) T → I (Cross, PGD-2) (Uni, BERT) All 83.4 66.8 79.3 64.2 71.0 49.4 33.3 22.8 8.74 $\times$ 0.99
(1-3) T → I (CLIP, PGD-2) (CLIP, BERT) All - - - - - - - - 460.7 $\times$ 52.0
Perturbation order ablation
(2-1) I → T (Cross, PGD-2) (Cross, BERT) All 84.4 68.7 80.5 65.0 74.8 51.2 36.3 24.6 8.79 $\times$ 1.00
(2-2) T → I → T (Cross, PGD-2) (Cross, BERT) All 83.9 67.3 79.8 64.0 74.6 53.1 38.1 25.9 13.97 $\times$ 1.59
(2-3) I → T → I (Cross, PGD-2) (Cross, BERT) All 82.3 65.7 79.9 62.7 69.3 49.1 37.4 25.1 11.10 $\times$ 1.26
Perturbation strength ablation
(3-1) T → I (Cross, PGD-2) EDA All 86.2 71.5 82.5 67.4 72.3 48.6 35.5 22.5 3.58 $\times$ 0.41
(3-2) T → I (Cross, PGD-10) EDA All 84.3 67.8 80.9 65.7 68.4 45.7 36.0 23.7 12.18 $\times$ 1.39
(3-3) T → I (Cross, PGD-10) (Cross, BERT) All 79.9 63.8 78.0 61.6 68.2 47.7 38.5 24.7 17.22 $\times$ 1.96
Table 3: Analysis of MAT’s strategies. Ablation study of CLIP trained on Flickr30k for image-text retrieval (ITR), comparing attack objectives (Cross- vs. Uni-modal), perturbation order (“Order”), and attack strength (PGD-2 vs. PGD-10, EDA vs. BERT).
@@
(ours) MAT+ I2T(Human) 71.0 $\uparrow$4.1 54.3 $\uparrow$4.5 35.6 $\uparrow$4.3 25.7 $\uparrow$4.7
@@
Datasets. We use the commonly used Flickr30k [plummer2015flickr30k] and the COCO [chen2015microsoft] datasets for ITR. While these datasets contain five captions per image, our baseline training uses 1:1 image-text pairs (taking the first annotated caption), reflecting the typical setup when fine-tuning with real-world data, such as data collected from the internet, where 1:1 pairings are more common. For VG, we use RefCOCO+ [kazemzadeh2014referitgame], and for IC, we use COCO. See Appendix A for details.
Evaluation metrics. We use the recall@k (R@k) for evaluating ITR, including both image-to-text and text-to-image retrieval. For VG we measure the accuracy in localizing the regions corresponding to the text descriptions, while IC evaluates the quality of generated captions using diverse metrics. Details are in Appendix A.
Adversarial attacks. We evaluate defense methods against the multimodal adversarial attack SGA [lu2023set], with perturbation constraints of $\epsilon_{I}=2/255$ ($L_{\infty}$-norm) for images and one token for text, following [zhang2022towards, lu2023set]. Results on unimodal attacks (PGD, BERT-Attack) and Co-Attack [zhang2022towards] are in Appendix B.
Baseline defense methods. Since there do not exist defense methods tailored for multimodal attacks, we compare our method with the unimodal (image-only) AT methods closest to our setting proposed for CLIP:
  • FARE [schlarmann2024robust]: An unsupervised (unimodal) adversarial fine-tuning scheme for CLIP, which focuses on obtaining a robust CLIP vision encoder. Since it is unsupervised, we can apply it to our ITR setting as-is.
  • TeCoA-ITR: Since the original TeCoA [mao2022understanding] is text-guided, we extend it to be used in ITR. While the original TeCoA fine-tunes the vision encoder with image classification loss, TeCoA-ITR fine-tunes all parameters using cross-modal objective of Eq. 8 to generate adversarial images.
Training details. We fine-tune the pre-trained CLIP-ViT-B/16, ALBEF-14M, and BLIP w/ ViT-B models using MAT. Adversarial images are generated via 2-step-PGD (perturbation size of 2/255 in $l_{\infty}$-norm), and adversarial texts using BERT-attack (1-token perturbation). Please see the detailed hyperparameters in Appendix A.4. Computational cost details are in Appendix A.5.
@@
    • $\circ$
      Intra-modal: EDA [wei2019eda] for basic word-level edits. LLM-based rewriting [fan2024improving] is shown in Appendix B.
    • $\circ$
      Cross-modal: Image-to-text (I2T) generation using diverse prompts with InternVL [chen2024internvl] (“I2T(div-Caps)”), and human-generated captions from the remaining 4 annotations of Flickr30k and COCO (“I2T(Human)”).
@@
    • $\circ$
      Intra-modal: RandAug [cubuk2020randaugment], randomly applying affine transformations and color distortions.
    • $\circ$
      Cross-modal: Stable Diffusion (SD) for text-to-image (T2I) [rombach2022high] and text-guided image-to-image (T-I2I) [meng2021sdedit].
@@ 6.1.1 Effectiveness of Multimodal Adversarial Training
Defending against multimodal attacks requires multimodal perturbations. With multimodal adversarial perturbations, MAT consistently achieves significantly greater robustness against multimodal attacks than the unimodal AT methods, FARE and TeCoA-ITR, which focus solely on image perturbations during AT. The improvements are substantial and consistent for CLIP on Flickr30k and COCO (Tab. 1), as well as ALBEF on both datasets (Tab. 2). These results highlight the necessity of defense strategies tailored for multimodal perturbations.
Extensive analysis towards efficient and effective multimodal defense. Solving the inner-maximization in Eq. 6 is non-trivial, since it requires updating both modalities and involves a high computational cost. We conducted ablation studies on key design factors—objective functions, perturbation order, and perturbation strength—to make our defense efficient and effective (Tab. 3).
  • Objective functions: Crossmodal loss (cosine similarity between image-text pairs) outperforms unimodal loss (similarity between orig. and adv. samples), with image-side objective being particularly critical since the image modality is more vulnerable than text (MAT vs. (1-1)/(1-2)). For text perturbations, optimizing the full CLIP loss in BERT-Attack is prohibitively slow ((1-3)); disrupting a single image-text pair suffices with much lower cost.
[…]
  • Perturbation strength: Image: Strength is controlled by PGD iterations. While FARE and TeCoA-ITR use a ten-steps PGD (PGD-10), MAT achieves comparable efficiency by using PGD-2, sufficient for multimodal robustness. Using PGD-10 further improves robustness, at the cost of clean accuracy and efficiency ((3-3)). Text: EDA (simple word edits) is an efficient alternative to BERT-Attack ((3-1), (3-2)), since single-token perturbations in the discrete space are largely covered by simple edits.
6.1.2 Effectiveness of one-to-many augmentations